World Wide Web Wares​ (4W, for​ short) is an online retailer of small kitchen appliances and utensils. The firm has been around for a few years and has created a nice market niche for itself. In​ fact, it actually turned a profit last​ year, albeit a fairly small one. After doing some basic research on the​ company, you've decided to take a closer look. You plan to use the​ price-to-sales ratio to value the​ stock, and you have collected​ P/S multiples on the following Internet retailer​ stocks:  ​(Click on the icon located on the​ top-right corner of the data table below in order to copy its contents into a​ spreadsheet.)

Company

​P/S Multiples

Amazing.com

4.7

Really Cooking.com

4.2

Fixtures​ & Appliances Online

3.9

Find the average​ P/S ratio for these three firms. Given that 4W is expected to generate $32 million in sales next​ year, and will have 5million shares of stock​ outstanding, use the average​ P/S ratio you computed above to put a value on​ 4W's stock.

The average​ P/S ratio for the 3 firms is ? ​(Round to two decimal​ places.)

The value of​ 4W's stock is ? (Round to the nearest​ cent.)
Answer #1
Company P/S Ratio
Amazing.com 4.7
Really Cooking.com 4.2
Fixtures & Appliances Online 3.9
Average 4.27

The average P/S ratio for the 3 firms is 4.27

The value of the firm = (P/S Multiple * Sales) / No.of Shares

Value of the firm = (4.27*32)/5 = $27.31 per share